Take a look at this new companies projected balance sheet. Analyze and describe the health of the company.

2020

2021

2022

Cash

$593,561

$674,051

$1,169,755

Accounts Receivable

$54,538

$84,079

$138,615

Inventory

Other Current Assets

$0

$0

Total Current Assets

$648,098

$758,130

$1,308,369

Long-Term Assets

Accumulated Depreciation

Total Long-Term Assets

Total Assets

$648,098

$758,130

$1,308,369

Accounts Payable

$11,437

$57,239

$88,618

Income Taxes Payable

$18,095

$5,556

$30,107

Sales Taxes Payable

$11,537

$19,403

$31,988

Short-Term Debt

$0

$0

$0

Prepaid Revenue

Total Current Liabilities

$41,069

$82,198

$150,713

Long-Term Debt

$20,000

$0

$0

Total Liabilities

$61,069

$82,198

$150,713

Paid-In Capital

$514,650

$514,650

$514,650

Retained Earnings

$72,379

$161,282

Earnings

$72,379

$88,903

$481,724

Total Owner's Equity

$587,029

$675,932

$1,157,656

Total Liabilities & Equity

$648,098

$758,130

$1,308,369
Answer #1

With the help of Financial Ratios, we can determine the health of the company:

1) Working Capital = Current Assets - Current Liabilities

For 2020, WC = 6,48,098- 41,069 = 6,07,029 $

For 2021, WC = 758130- 82198 = 6,75,932$

For 2022, WC = 1308369 - 150713 = 11,57,656$

The company has good working capital which means that the company will be able to pay the current obligations (bills, payroll, etc)

2) Current Ratio = Current Assets/ Current Liabilities

For 2020, CR = 648098/ 41069 = 15.78

For 2021, CR = 758130/ 82198 = 9.22

For 2022, CR = 1308369/ 150713 = 8.68

The current ratio, although decreasing over a period of time is quite high, which means the health of the company is okay

3) Debt to Equity Ratio = Total Liabilities / Total Shareholder's Equity

For 2020 , Debt to Equity Ratio = 61069/ 587029 = 0.10

For 2021, Debt to Equity Ratio = 82198/ 675932 = 0.12

For 2022, Debt to Equity Ratio = 150713/ 1157656 = 0.13

The Debt to Equity Ratio should always be low, so the health of the company is good. Although, the low ratio indicates that company is not taking advantage of increased profits, that is something which can be looked in to.

From these ratios, the health of the company looks good however the ratios Inventory Turnover Ratio, Receivable turnover ratio ,Payable Turnover Ratio and Asset Turnover Ratio gives a better indication of the health of the company, in this case it cannot be calculated since the value of Inventory and Net sales has not been specified directly.